OpenStudy (anonymous):

During an intramural basketball game, Team A scored 18 fewer points than Team B. Together, both teams scored a total of 148 points. How many points did Team A score during the game? a. 66 points b. 83 points c. 74 points d. 65 points

OpenStudy (anonymous):

d 65 points

OpenStudy (anonymous):

d. 65 points

OpenStudy (anonymous):

It's d. Do you want justification?

OpenStudy (anonymous):

if team A scored x points then team B scored x + 18 points x + x + 18 = 148 2x = 148-18 x = 130/2 = 65

OpenStudy (anonymous):

Yeah! already did it (got it right) and you cleared my head on weather it was right! Thanks!

OpenStudy (anonymous):

whether! =P